a man is 30 years old while his daughter is 4 years old. In how many years' time will the daughter be half the age of her father?
Answer by Cromlix(4381)

a man is 30 years old while his daughter is 4 years old.
In how many years' time will the daughter be half the age of her father?
30 + x = 2(4 + x)
30 + x = 8 + 2x
30 - 8 = 2x - x
x = 22
In 22 years time the daughter will
be half the father's age.
Hope this helps.
:-)